### v1

@@ 2.4.1 Time hop ($t_{h}$) and Learning rate ($\eta$)
As it can be seen, $t_{h}$ values between $1.0$ s. and $2.0$ s. provide the least amount of SDR variance. With this information, another evaluation was carried out to find the best combination of $t_{h}$ and $\eta$ values. To do this, for every combination, $10$ runs were carried out (obtaining the value of $\theta_{corr}$ that maximizes $Q$), with varying amounts of location errors ($\theta_{est}$). A ‘good run’ is defined as one in which the root-mean-square of the segment of the last $3/4$ of its $\theta_{corr}$ values is in the ±$\pm 5^{\circ}$ range. This segment was chosen since the first $1/4$ of the run is where the optimization process is expected to be still converging into the correct DOA, and should not affect the result of establishing it as a ‘good run’. Since there was a considerable amount of evaluated combinations ($t_{h}=[1.0,1.5,2.0]$, $\eta=[0.2,0.25,0.3,0.35,0.4,0.5,0.6]$, $\theta_{est}=[1,5,10,15,20,25]$), only the top $10$ combinations (with the highest amount of average ‘good runs’ throughout all $\theta_{est}$ values) are reported in Table 5.
@@
$\theta_{est}$ range $t_{h}$ (s) $\eta$ good runs
$[1^{\circ},5^{\circ}]$ $2.0$ $0.20$ $8.0$
$[1^{\circ},5^{\circ}]$ $1.0$ $0.20$ $7.0$
$[1^{\circ},5^{\circ}]$ $1.5$ $0.20$ $7.0$
$[10^{\circ},15^{\circ}]$ $1.5$ $0.30$ $6.0$
$[10^{\circ},15^{\circ}]$ $1.5$ $0.20$ $5.0$
$[10^{\circ},15^{\circ}]$ $1.5$ $0.35$ $5.0$
$[20^{\circ},25^{\circ}]$ $0.5$ $0.25$ $2.5$
$[20^{\circ},25^{\circ}]$ $1.0$ $0.25$ $2.0$
$[20^{\circ},25^{\circ}]$ $1.5$ $0.30$ $1.5$
@@ 2.4.2 New Optimization Mechanism with DOA Merger
As a reminder, the optimization task aims to modify one input parameter ($\theta_{corr}$) to optimize one observable variable ($Q$). The insight of this new optimization mechanism is as follows: at any given moment, the correction process can recognize that it has overshot the correct DOA by remembering the $\theta_{corr}$ that has provided the highest $Q$ up until to that point ($Q_{best}$), and ‘reset’ itself if necessary. Thus, if none of the recent $\theta_{corr}$ values have not provided a higher $Q$ than $Q_{best}$, the new optimization mechanism reverts back to the $\theta_{corr}$ value that provided it ($\theta_{best}$) and re-starts the optimization process from there. This new mechanism is detailed in Algorithm 1, where all variables are initialized at 00, unless stated otherwise.
1:$t_{h}$: time hop; $\eta$: learning rate
2:$\beta_{m}$: momentum factor ($0.9$); $\beta_{v}$: variance factor ($0.999$)
@@
4:Initialize: $\theta_{c}\leftarrow\theta_{est}$triangleright from soundloc agent
@@
[…]
@@
16:                  $\theta_{c}\leftarrow\theta_{best}$triangleright reset to best DOA
@@
25:     $\theta_{c}\leftarrow merge\_theta(\theta_{c},\theta_{est})$triangleright from soundloc agent
26:     $\theta_{c}\to\theta_{corr}$triangleright to speech enhancement module
@@
29:     $Q_{c}\leftarrow Q$triangleright from onlinesqa agent
@@
There are two functions used in Algorithm 1. The $quality\_norm$ function normalizes the value of a quality value ($Q$) such that its value is aimed to be minimized (as the Adam implementation expects it to be). This is carried out by subtracting $Q$ from its maximum value ($Q_{MAX}$): $quality\_norm(Q):=Q_{MAX}-Q$. In the case of the SDR metric, $Q_{MAX}=100$.
As for the $merge\_theta$ function, it aims to merge the DOA estimation that is continuously provided by the soundloc agent ($\theta_{est}$) and the corrected DOA ($\theta_{corr}$). It does so by implementing the concept of ‘boredom’ and ‘excitement’, based on the variability of $\theta_{est}$ and $\theta_{corr}$. Basically, the weighted standard deviation of the linearly detrended $\theta_{est}$ history ($\sigma_{est_{hist}}$) establishes if there has been a recent important change in the output of the soundloc agent. If not, the system is deemed ‘bored’, and continues with the normal execution of Algorithm 1, only outputting $\theta_{corr}$. If there has been an important recent change in $\sigma_{est_{hist}}$, the system is deemed ‘excited’ and, based on the ratio between the recent variability of $\theta_{est}$ and $\theta_{corr}$, a new $\theta_{corr}$ is calculated, such that the $\theta$ with more variability is given more preference. Algorithm 2 details this process.
To evaluate the impact of this new optimization mechanism, $10$ runs were carried out with $\eta$ set at $0.3$ and $t_{h}$ set at $1.5$ s. (as recommended by the results presented in Section 2.4.1), with different values of $\theta_{est}$ ($[1,5,10,15,20,25]$). Each of those sets of runs were ran with both the original optimization mechanism, as well as with the proposed new optimization mechanism. The average RMS ($\theta_{error}$), their standard deviation ($\theta_{var}$), as well as the number of good runs per each $[\eta,t_{h},\theta_{est}]$ combination are presented in Table 7.
@@
$\theta_{error}$ $\theta_{var}$ good runs
@@
$1^{\circ}$ $6.72$ $\mathbf{4.96}$ $3.02$ $\mathbf{2.20}$ $5$ $\mathbf{6}$
$5^{\circ}$ $4.85$ $\mathbf{3.77}$ $2.51$ $\mathbf{0.54}$ $7$ $\mathbf{10}$
$10^{\circ}$ $4.93$ $\mathbf{4.02}$ $2.41$ $\mathbf{0.67}$ $6$ $\mathbf{9}$
$15^{\circ}$ $6.19$ $\mathbf{5.99}$ $3.63$ $\mathbf{2.67}$ $4$ $\mathbf{5}$
$20^{\circ}$ $9.28$ $\mathbf{7.03}$ $6.27$ $\mathbf{3.65}$ $2$ $\mathbf{4}$
$25^{\circ}$ $\mathbf{11.31}$ $12.82$ $\mathbf{7.33}$ $7.34$ 00 00
@@
1:$\theta_{hist}$: history of $\theta_{c}$ from Alg. 1
2:$\theta_{est_{hist}}$: history of $\theta_{est}$, detrended linearly $\triangleright$ from soundloc agent
3:$i_{bored}$: current number of ‘bored’ windows
4:$N_{bored}$: maximum number of ‘bored’ windows
5:$\sigma_{bored}$: variance threshold to establish ‘boredom’
@@
7:$w_{est_{hist}}=\frac{1}{1+e^{-\left(\left[1...N_{est_{hist}}\right]-\frac{N_{est_{hist}}}{2}\right)}}$
8:$\overline{\theta_{est_{hist}}}\leftarrow\frac{\sum_{i}\theta_{est_{hist}}[i]}{N_{est_{hist}}}$
9:$\sigma_{est_{hist}}\leftarrow\sqrt{\frac{\sum_{i}(\theta_{est_{hist}}[i]-\overline{\theta_{est_{hist}}})*(\theta_{est_{hist}}[i]w_{est_{hist}}[i]-\overline{\theta_{est_{hist}}})}{N_{est_{hist}}-1}}$
@@
12:         $i_{bored}\leftarrow 0$triangleright Excited, start using $\theta_{est}$
@@
[…]
@@
25:         $i_{bored}\leftarrow 0$triangleright Keep being excited
@@
27:         $i_{bored}\leftarrow i_{bored}+1$triangleright Starting to get bored
@@
Algorithm 2 $merge\_theta$ function.
To further illustrate these benefits, Figures 8 and 9 show the original and the new optimization mechanism’s behavior (respectively) with $\theta_{est}=10^{\circ}$. It can be seen how the variance is reduced at the end of the runs, with a more consistent stream of $\theta_{corr}$ values within the ±$\pm 5^{\circ}$ range.
@@ 2.5 Frequency Selection based on Source Type
In Figure 14, the output of the soundloc agent is shown when it masked the peak frequencies of the urban sound, which are $[999,1229,1485,1630,2005,2478,2844,2985,3257,3486,3670,3976]$ Hz. These were extracted from the log-frequency domain, calculated from the urban sound recording that was inserted into the input audio mixture.

### v2

@@ 2.4.1 Time hop ($t_{h}$) and Learning rate ($\eta$)
As it can be seen, $t_{h}$ values between $1.0$ s. and $2.0$ s. provide the least amount of SDR variance. With this information, another evaluation was carried out to find the best combination of $t_{h}$ and $\eta$ values. To do this, for every combination, $10$ runs were carried out (obtaining the value of $\theta_{corr}$ that maximizes $Q$), with varying amounts of location errors ($\theta_{est}$). A ‘good run’ is defined as one in which the root-mean-square of the segment of the last $3/4$ of its $\theta_{corr}$ values is in the $\pm 5^{\circ}$ range. This segment was chosen since the first $1/4$ of the run is where the optimization process is expected to be still converging into the correct DOA, and should not affect the result of establishing it as a ‘good run’. Since there was a considerable amount of evaluated combinations ($t_{h}=[1.0,1.5,2.0]$, $\eta=[0.2,0.25,0.3,0.35,0.4,0.5,0.6]$, $\theta_{est}=[1,5,10,15,20,25]$), only the top $10$ combinations (with the highest amount of average ‘good runs’ throughout all $\theta_{est}$ values) are reported in Table 5.
@@
$\theta_{est}$ range $t_{h}$ (s) $\eta$ # good runs
$1^{\circ},5^{\circ}]$ $2.0$ $0.20$ $8.0$
$1^{\circ},5^{\circ}]$ $1.0$ $0.20$ $7.0$
$1^{\circ},5^{\circ}]$ $1.5$ $0.20$ $7.0$
$10^{\circ},15^{\circ}]$ $1.5$ $0.30$ $6.0$
$10^{\circ},15^{\circ}]$ $1.5$ $0.20$ $5.0$
$10^{\circ},15^{\circ}]$ $1.5$ $0.35$ $5.0$
$20^{\circ},25^{\circ}]$ $0.5$ $0.25$ $2.5$
$20^{\circ},25^{\circ}]$ $1.0$ $0.25$ $2.0$
$20^{\circ},25^{\circ}]$ $1.5$ $0.30$ $1.5$
@@ 2.4.2 New Optimization Mechanism with DOA Merger
As a reminder, the optimization task aims to modify one input parameter ($\theta_{corr}$) to optimize one observable variable ($Q$). The insight of this new optimization mechanism is as follows: at any given moment, the correction process can recognize that it has overshot the correct DOA by remembering the $\theta_{corr}$ that has provided the highest $Q$ up until to that point ($Q_{best}$), and ‘reset’ itself if necessary. Thus, if none of the recent $\theta_{corr}$ values have not provided a higher $Q$ than $Q_{best}$, the new optimization mechanism reverts back to the $\theta_{corr}$ value that provided it ($\theta_{best}$) and re-starts the optimization process from there. This new mechanism is detailed in Algorithm 1, where all variables are initialized at 0, unless stated otherwise.
@@
[…]
3:Initialize: $\theta_{hist}\leftarrow[0,0,\dots,0]$; $i_{nocormax}\leftarrow 10$; $\theta_{best}\leftarrow\mathit{None}$; $Q_{best}\leftarrow Q_{MAX}$
4:Initialize: $\theta_{c}\leftarrow\theta_{est}$ $\triangleright$ from soundloc agent
@@
9:  if $i_{calc}\geq\mathit{length}(\theta_{hist})$ then
@@
16:      $\theta_{c}\leftarrow\theta_{best}$ $\triangleright$ reset to best DOA
@@
25:  $\theta_{c}\leftarrow merge\_theta(\theta_{c},\theta_{est})$ $\triangleright$ from soundloc agent
26:  $\theta_{c}\to\theta_{corr}$ $\triangleright$ to speech enhancement module
@@
29:  $Q_{c}\leftarrow Q$ $\triangleright$ from onlinesqa agent
@@
There are two functions used in Algorithm 1. The quality_normquality\_norm function normalizes the value of a quality value ($Q$) such that its value is aimed to be minimized (as the Adam implementation expects it to be). This is carried out by subtracting $Q$ from its maximum value ($Q_{MAX}$): $quality\_norm(Q):=Q_{MAX}-Q$. In the case of the SDR metric, $Q_{MAX}=100$.
As for the merge_thetamerge\_theta function, it aims to merge the DOA estimation that is continuously provided by the soundloc agent ($\theta_{est}$) and the corrected DOA ($\theta_{corr}$). It does so by implementing the concept of ‘boredom’ and ‘excitement’, based on the variability of $\theta_{est}$ and $\theta_{corr}$. Basically, the weighted standard deviation of the linearly detrended $\theta_{est}$ history ($\sigma_{est_{hist}}$) establishes if there has been a recent important change in the output of the soundloc agent. If not, the system is deemed ‘bored’, and continues with the normal execution of Algorithm 1, only outputting $\theta_{corr}$. If there has been an important recent change in $\sigma_{est_{hist}}$, the system is deemed ‘excited’ and, based on the ratio between the recent variability of $\theta_{est}$ and $\theta_{corr}$, a new $\theta_{corr}$ is calculated, such that the $\theta$ with more variability is given more preference. Algorithm 2 details this process.
To evaluate the impact of this new optimization mechanism, $10$ runs were carried out with $\eta$ set at $0.3$ and $t_{h}$ set at $1.5$ s. (as recommended by the results presented in Section 2.4.1), with different values of $\theta_{est}$ ($[1,5,10,15,20,25]$). Each of those sets of runs were ran with both the original optimization mechanism, as well as with the proposed new optimization mechanism. The average RMS ($\theta_{error}$), their standard deviation ($\theta_{var}$), as well as the number of good runs per each $\eta,t_{h},\theta_{est}]$ combination are presented in Table 7.
@@
$\theta_{error}$ $\theta_{var}$ # good runs
@@
$1^{\circ}$ $6.72$ $\mathbf{4.96}$ $3.02$ $\mathbf{2.20}$ 55 $\mathbf{6}$
$5^{\circ}$ $4.85$ $\mathbf{3.77}$ $2.51$ $\mathbf{0.54}$ 77 $\mathbf{10}$
$10^{\circ}$ $4.93$ $\mathbf{4.02}$ $2.41$ $\mathbf{0.67}$ 66 $\mathbf{9}$
$15^{\circ}$ $6.19$ $\mathbf{5.99}$ $3.63$ $\mathbf{2.67}$ 44 $\mathbf{5}$
$20^{\circ}$ $9.28$ $\mathbf{7.03}$ $6.27$ $\mathbf{3.65}$ 22 $\mathbf{4}$
$25^{\circ}$ $\mathbf{11.31}$ $12.82$ $\mathbf{7.33}$ $7.34$ 0 0
@@
[…]
@@
12:   $i_{bored}\leftarrow 0$ $\triangleright$ Excited, start using $\theta_{est}$
@@
19:  $N_{hist}\leftarrow\mathit{length}(\theta_{hist})$
@@
25:   $i_{bored}\leftarrow 0$ $\triangleright$ Keep being excited
@@
27:   $i_{bored}\leftarrow i_{bored}+1$ $\triangleright$ Starting to get bored
@@
Algorithm 2 merge_thetamerge\_theta function.
To further illustrate these benefits, Figures 8 and 9 show the original and the new optimization mechanism’s behavior (respectively) with $\theta_{est}=10^{\circ}$. It can be seen how the variance is reduced at the end of the runs, with a more consistent stream of $\theta_{corr}$ values within the $\pm 5^{\circ}$ range.
@@ 2.5 Frequency Selection based on Source Type
In Figure 14, the output of the soundloc agent is shown when it masked the peak frequencies of the urban sound, which are [999,1229,1485,1630,2005,2478,2844,2985,3257,3486,3670,3976][999,1229,1485,1630,2005,2478,2844,2985,3257,3486,3670,3976] Hz. These were extracted from the log-frequency domain, calculated from the urban sound recording that was inserted into the input audio mixture.